OpenStudy (anonymous):

Given tanx=1/2 and sinx < 0, find sin2x,cos2x, and tan2x. I found tan2x but i need help with sin2x and cos2x

OpenStudy (anonymous):

Clearly, we have \[ \sin(x)=\frac{\cos(x)}{2} \] Thus, \[ \sin^2(x)+\cos^2(x)=\frac{\cos^2(x)}{4}+\cos^2(x) \] \[ 1= \frac{5}{4}\cos^2(x) \] Since \[ \sin(x) <0\]therefore \[ \cos(x) <0 \] From the equation above, we derive at \[ \cos(x)=\frac{-2}{\sqrt{5}} \] and \[ \sin(x)=\frac{-1}{\sqrt{5}} \] Moreover, using double angle formula, we have \[ \sin(2x)=2\sin(x)\cos(x) \\ \cos(2x)=\cos^2(x)-\sin^2(x) \]

OpenStudy (anonymous):

 can u show me how to draw it on a quadrant plane? i know its in the 3rd one but i'm confused on how to draw it

OpenStudy (anonymous):



OpenStudy (anonymous):

you are right, it is in the third quadrant. The moral lesson is that you should be more comfortable with your answer, believe in your instinct.

OpenStudy (anonymous):

i understand. thx ^_^